40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 14 18:45:43 2022, Anchal, Update, SUS, LO2 lower magnets are stuck in coil, won't come off 
    Reply  Tue Jun 14 19:34:06 2022, yuta, Update, SUS, Resonant frequency identification from the free swing test 7x
       Reply  Thu Jun 16 18:30:07 2022, Anchal, Update, SUS, New diagonalized input matrices applied SUSInpMatDiag_MC1_MC2_MC3_BS_IMTX_ETMX_ETMY_AS1_AS4_SR2_PR2_PR3.pdf
          Reply  Tue Jun 21 08:36:50 2022, Anchal, Update, SUS, Diagonalized input matrices for LO1, freeSwing on ITMY and ITMX SUSInpMatDiag_LO1.pdf
    Reply  Wed Jun 15 15:07:07 2022, Koji, Update, SUS, LO2 SUS stuck fixed Screenshot_2022-06-15_15-00-22.pngScreenshot_2022-06-15_15-00-40.png
Message ID: 16931     Entry time: Tue Jun 21 08:36:50 2022     In reply to: 16925
Author: Anchal 
Type: Update 
Category: SUS 
Subject: Diagonalized input matrices for LO1, freeSwing on ITMY and ITMX 

Over the weekend, I ran freeSwing test with sequential kicks in specific DOFs for LO1, ITMY, ITMX. LO1 results were successfully used to diagonalize LO1 input matrix. There are some issues for ITMY, ITMX still. I could not run LO2 test.


LO1

The free swing test ran successfully, resonant frequencies for different DOFs was extracted, and new input matrix was calculated. The new matrix was only slightly different from before and it worked fine with existing damping loops. The observed resonance frequencies were different from previous values by POS: -6 mHz, PIT : -3 mHz, YAW: -9 mHz, SIDE: -2 mHz. Attached are the diagonalization result.


ITMX

The peculiarity of ITMX remained even after the second free swing test. The calculated input matrix is very different from existing one with sign flips across PIT and POS rows. I found that our LR osem is always bright in ITMX at the current alignment position. I see that LR osem comes in range when C1:SUS-ITMX_PIT_COMM is raised above 0.5. Maybe we should run this test when we know for sure ITMX is in correct position.


ITMY

In ITMY on the other hand, I found that SIDE OSEM was completely bright. This happened during the YAW kick to ITMY. We'll need to reduce kick amplitudes for ITMY and redo this test.


LO2

For LO2, I could not initiate the test. On reducing the alignment offsets for LO2 (so that it doesn't get stuck in the fre swing test), the damping loops were not working. This is a clear evidence also that input matrix is different for different positions of the optic. We need to think about some other strategy to do this test, maybe see if ideal input amtrix works at no offsets and use that to damp during the test.


Attachment 1: SUSInpMatDiag_LO1.pdf  75 kB  Uploaded Tue Jun 21 09:45:01 2022  | Hide | Hide all
SUSInpMatDiag_LO1.pdf
ELOG V3.1.3-